# Less LLM, More Documents: Searching for Improved RAG

### v2

@@ 6 Conclusion
In this paper, we asked whether scaling the retrieval corpus can often substitute for scaling the generator in RAG, and how corpus size interacts with model size under a fixed evidence budget. Using controlled evaluations on NQ, TriviaQA, and WebQ with standardized prompting and context formatting, we characterize the corpus–generator trade-off while holding the presented evidence constant.
Our results show that corpus scaling can often offset model downsizing. Across datasets, enlarging the corpus is a reliable lever: smaller or mid-sized generators paired with larger corpora frequently surpass larger models under the same evidence budget. In several settings, moving up corpus tiers closed the gap of one to two model-size tiers, demonstrating that “more documents” can often substitute for “more parameters” when inference resources are constrained.
[…]
Practically, when resources constrain generator size, it is often better to expand the corpus. Pairing mid-sized generators with larger corpora effectively converts coverage into end-task gains, whereas very large models offer little additional benefit, and very small ones require steep corpus expansions. Although we mainly focus on the Qwen3 family due to the lack of other open-source LLM series with homogeneous, wide-ranging variants, we hope to extend the analysis to additional model families once they become available. Notably, mid-sized models sometimes exploit retrieved context more efficiently than the largest models, consistent with our utilization analysis. Tracking gold-answer coverage and the Utilization Ratio provides practical diagnostics to guide budgeting between retriever and generator. In short: Less LLM, More Documents.

### v3

@@ 6 Conclusion
In this paper, we asked whether scaling the retrieval corpus can substitute for scaling the generator in RAG, and how corpus size interacts with model size under a fixed evidence budget. Using controlled evaluations on NQ, TriviaQA, and WebQ with standardized prompting and context formatting, we characterize a corpus–generator trade-off while holding the presented evidence constant.
Our results indicate that, in this open-domain setting, corpus scaling can often partially offset model downsizing. Across datasets, enlarging the corpus is a reliable lever: smaller or mid-sized generators paired with larger corpora frequently approach, and sometimes exceed, the accuracy of larger models under the same evidence budget. In several settings, moving up corpus tiers closes the gap of one to two model-size tiers, suggesting that increasing corpus scale can in many cases yield gains comparable to moving to a larger generator tier.
A consistent mechanism explains these gains: improvements are primarily driven by increased coverage of answer-bearing evidence rather than changes in utilization efficiency. As the corpus grows, the likelihood that retrieved passages contain the gold answer increases consistently, whereas the model’s context-utilization ratio remains roughly stable across shard counts and model sizes; we observe the same qualitative corpus-scaling trends under an auxiliary retriever built with a different encoder, suggesting the main findings are not specific to a single embedding model. Thus, scaling mainly works by raising the hit rate of relevant evidence, instead of altering how effectively models exploit the provided context. At the same time, both the paired bootstrap differences between adjacent shard counts and the per-shard CB increments $\Delta_{n}$ show that marginal gains rapidly decay and become indistinguishable from zero beyond roughly a $5\times$ corpus increase, indicating clear diminishing returns at larger scales.
From a practical perspective, our findings suggest that retrieval-side scaling is an important design axis when additional corpus is available and retrieval costs are acceptable under deployment constraints. Under our evidence budget, a simple serving-time FLOPs proxy indicates an orders-of-magnitude compute gap between single-shard retrieval and generation, suggesting that overall arithmetic compute is typically dominated by the generator (notwithstanding retrieval’s memory/IO and indexing overheads). In our experiments, mid-sized generators tend to convert improved evidence coverage into end-task gains more effectively than very small models (which require steep corpus expansions) and very large models (which yield smaller marginal gains). While we mainly focus on the Qwen3 family due to the lack of other open-source LLM series with homogeneous, wide-ranging variants, we hope to extend this analysis to additional model and retriever families, as well as more domain-specific and siloed corpora, in future work. Finally, tracking gold-answer coverage together with the Utilization Ratio provides practical diagnostics for when increasing retrieval scale is likely to translate into downstream improvements; when corpus expansion is limited (e.g., enterprise or specialized domains), the same evidence-centric view suggests a coverage-first strategy via domain-focused curation and retrieval quality improvements, rather than adding low-yield documents.